65 SCARLET AND GRAY BOOKLETS

\underline{BRENDA P. WINNEWISSER}, Department of Physics, The Ohio State University, Columbus, Ohio 43210.

Delving into the archives, and drawing on the experiences of various spectroscopists, several questions will be addressed, including: What has changed since the first meeting? What has not changed since the first meeting? How did this meeting spawn not only one, but two similar meetings in Europe? How has international participation evolved over the years? How has female participation evolved? And why do we keep coming back?